# Modeling Reactions on the Solid-Liquid Interface With Next Generation Extended Lagrangian Quantum-Based Molecular Dynamics

### v1

@@ 2 Introduction
In the new generation of extended Lagrangian first-principles molecular dynamics (XL-BOMD) methods developed by Niklasson et al. 32, 33, 34, 35, 36, 37, 38, 29, both the nuclear and extended electronic degrees of freedom are evolved. The electronic evolution follows a similar approach as in the Car-Parrinello QMD method39, 37. However, in contrast to Car-Parrinello, only a single full diagonalization of the electronic Hamiltonian is needed at each time step. Moreover, XL-BOMD avoids system-dependent electron mass parameters and allows longer integration time steps of the same order as in regular BOMD 37, 29. The error of the sampled Born-Oppenheimer potential energy surface is of the fourth order in the integration time step with a small prefactor. This accuracy can be increased to the eighth order if necessary 40. In XL-BOMD the electronic degrees of freedom evolve through a harmonic oscillator that closely follows the exact ground state during a QMD simulation. In its most recent incarnation, this oscillator includes an inner-product kernel which acts as a preconditioner 37, 38, 29, and can be efficiently approximated with a preconditioned low-rank Krylov subspace expansion 41.
@@ 3.2 Extended Lagrangian Born-Oppenheimer MD
Extended-Lagrangian Born-Oppenheimer MD guarantees a smooth evolution of the charge which appears as an auxiliary field variable in the equations of motion. XL-BOMD also allows us to use a relatively high integration time step while maintaining system stability. We will briefly present the XL-BOMD method as explained elsewhere 37, 29. Regardless of the underlying level of theory, in regular Born-Oppenheimer molecular dynamics, the potential energy surface, ${U}_{BO}(\textbf{R})$, with ${\bf R}=\{R_{i}\}_{i=1}^{N}$ being the nuclear coordinates, is determined through an iterative self-consistent optimization of a non-linear energy function of the electron density or charge distribution. In XL-BOMD, $U_{BO}({\bf R})$ is replaced by a shadow, or approximate, potential energy surface, ${\cal U}_{BO}(\textbf{R},n)$, which is linear with respect to some electron density $n(\textbf{r})$, where r are the electronic coordinates. The latter is obtained by performing a linear expansion of $U_{BO}({\bf R})$ around an approximate electronic ground state density, $\rho_{\rm min}(\textbf{r})$. Due to this linearization, the optimized density of the ground state, $\varrho[n(\textbf{r})]$ is given directly through a single diagonalization, avoiding expensive iterations. The density in XL-BOMD, $n(\textbf{r})$, is then included as a dynamic field variable that evolves through a harmonic oscillator that is centered around the exact density of the ground state, or the best approximation available to the exact ground state, that is, $\varrho[n(\textbf{r})]$. In this way ${\cal U}_{BO}(\textbf{R},n(\textbf{r}))$ closely follows the exact surface of the potential energy with an error that scales quadratically as ${\cal O}(||\varrho[n(\textbf{r})]-n(\textbf{r})||^{2})$ 37.
@@
$M_{i}\ddot{R}_{i}$ $=$ $-\left.\frac{\partial{\cal U}_{BO}(\mathbf{R},\textbf{q}[\textbf{n}])}{\partial R_{i}}\right|_{\textbf{n}},$
@@ 4 Results and Discussion
The simulations were run for 10 ps using a 0.1 fs time step and periodic boundary conditions. A Langevin thermostat was used to maintain a constant temperature of 300 K throughout the simulation 63 with a friction coefficient of $\gamma=0.01$ fs-1. More accurate methods could be used for larger friction values and may be explored in future work 64. Since these are potentially reactive systems, an electronic temperature of 0.025 [eV] / 8.617E-5 [eV/K] $\simeq$ 290 K was used to “smear” the Fermi distribution and improve the stability of the QMD simulations. Rank-N updates to the kernel were performed with the maximum rank set to 4.

### v2

@@ 2 Introduction
In the new generation of extended Lagrangian first-principles molecular dynamics (XL-BOMD) methods developed by Niklasson et al. 32, 33, 34, 35, 36, 37, 38, 29, 39, both the nuclear and extended electronic degrees of freedom are evolved. The electronic evolution follows a similar approach as in the Car-Parrinello QMD method40, 37. However, in contrast to Car-Parrinello, only a single full diagonalization of the electronic Hamiltonian is needed at each time step. Moreover, XL-BOMD avoids system-dependent electron mass parameters and allows longer integration time steps of the same order as in regular BOMD 37, 29. The error of the sampled Born-Oppenheimer potential energy surface is of the fourth order in the integration time step with a small prefactor. This accuracy can be increased to the eighth order if necessary 39. In XL-BOMD the electronic degrees of freedom evolve through a harmonic oscillator that closely follows the exact ground state during a QMD simulation. In its most recent incarnation, this oscillator includes an inner-product kernel which acts as a preconditioner 37, 38, 29, and can be efficiently approximated with a preconditioned low-rank Krylov subspace expansion 41.
@@ 3.2 Extended Lagrangian Born-Oppenheimer MD
Extended-Lagrangian Born-Oppenheimer MD guarantees a smooth evolution of the charge which appears as a dynamical field variable in the equations of motion. XL-BOMD also allows us to use a relatively long integration time step, as in regular Born-Oppenheimer MD, while maintaining system stability. We will briefly present the XL-BOMD method as explained elsewhere 37, 29. Regardless of the underlying level of theory, in regular Born-Oppenheimer molecular dynamics, the potential energy surface, ${U}_{\rm BO}(\textbf{R})$, with ${\bf R}=\{R_{i}\}_{i=1}^{N}$ being the nuclear coordinates, is determined through an iterative self-consistent optimization of a non-linear energy function of the electron density or charge distribution. In XL-BOMD, $U_{\rm BO}({\bf R})$ is replaced by a shadow, or approximate, potential energy surface, ${\cal U}_{\rm BO}(\textbf{R},n)$, which is based on the relaxed ground state of an approximate shadow energy functional that is based on a linearization of the regular energy density functional around an approximate ground state density $n(\textbf{r})$, where r are the electronic coordinates. Due to this linearization, the optimized density of the ground state, $\varrho[n(\textbf{r})]$ is given directly through a single diagonalization, avoiding expensive iterations. The density in XL-BOMD, $n(\textbf{r})$, is then included as a dynamic field variable that evolves through a harmonic oscillator that is centered around the exact density of the ground state, or the best approximation available to the exact ground state, that is, $\varrho[n(\textbf{r})]$. In this way ${\cal U}_{\rm BO}(\textbf{R},n(\textbf{r}))$ closely follows the exact surface of the potential energy with an error that scales quadratically as ${\cal O}(||\varrho[n(\textbf{r})]-n(\textbf{r})||^{2})$ 37. The accuracy of the shadow potential can be further improved by a first-level update 39 which has been used in these simulations.
@@
$M_{i}\ddot{R}_{i}$ $=$ $-\left.\frac{\partial{\cal U}_{\rm BO}(\mathbf{R},\textbf{q}[\textbf{n}])}{\partial R_{i}}\right|_{\textbf{n}},$
@@ 4 Results and Discussion
The simulations were run for 10 ps using a 0.1 fs time step and periodic boundary conditions. A Langevin thermostat was used to maintain a constant temperature of 300 K throughout the simulation 63 with a friction coefficient of $\gamma=0.01$ fs-1. More accurate methods could be used for larger friction values and may be explored in future work 64. Since these are potentially reactive systems, an electronic temperature of 0.025 [eV] / 8.617E-5 [eV/K] $\simeq$ 290 K was used to “smear” the Fermi distribution and improve the stability of the QMD simulations. Rank-N updates to the kernel were performed with the maximum rank set to 4. Simulations were run on the Los Alamos National Laboratory Institutional Computing Chicoma supercomputing cluster, using a Nvidia A100 GPU for acceleration. Each time step took $\sim$0.61 seconds wall-clock time using the GPMDK version that was available at the time production simulations were run.